Click here to read the eBook: Asset Management Ratios
Problem Walk-Through

DSO AND ACCOUNTS RECEIVABLE

Ingraham Inc. currently has $365,000 in accounts receivable, and its days sales outstanding (DSO) is 53 days. It wants to reduce its DSO to 35 days by pressuring more of its customers to pay their bills on time. If this policy is adopted, the company's average sales will fall by 20%. What will be the level of accounts receivable following the change? Assume a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.

$  
Answer #1

Days sales outstanding DSO=Receivables/Sales
Hence,
Receivables/Sales*365=53
=>Sales=Receivables*365/53=365000*365/53=2513679.245

New accounts receivable=35/365*365000*365/53*(1-20%)=192830.1887